Compound interest of 2 year is rs 156 and 3 year is rs 254. What is rate of interest? 

Grade:8

3 Answers

Sourabh Singh IIT Patna
7 years ago
hii

Difference between compound interest for 2 years and 3 years can be found out using the formulas P(1 + R/100)2 and P(1 + R/100)3 .

use this subtract them and you will get the answer where
P(1 + R/100)2 = 156 +P

P(1 + R/100)3 = 254 + P
sraddhansu sekhar dash
6 years ago
p{(1+r/100)^3 - 1}=254p{(1+r/100)^2 - 1}=156divide this 2 equation,(1+r/100)^3 - 1 / (1+r/100)^2 - 1=254/156=127/78let 1+r/100=xso, x^3-1/x^2-1=127/78put the formula of x^3 - 1^3 and x^2 - 1^2then,x-y will be cancelled..then, x^2+x+1 / x+1 = 127/7878x^2+78x+78=127x+12778x^2-49x-49=078x^2-91x+42x-49=013x(6x-7)+7(6x-7)=06x-7=0,x=7/6now x=1+r/100=7/6r/100=7/6-1=1/6r=(1/6)*100=50/3=16(2/3)...this is ur answer!!😊
sraddhansu sekhar dash
6 years ago
p{(1+r/100)^3 - 1}=254p{(1+r/100)^2 - 1}=156divide this 2 equation,(1+r/100)^3 - 1 / (1+r/100)^2 - 1=254/156=127/78let 1+r/100=xso, x^3-1/x^2-1=127/78put the formula of x^3 - 1^3 and x^2 - 1^2then,x-y will be cancelled..then, x^2+x+1 / x+1 = 127/7878x^2+78x+78=127x+12778x^2-49x-49=078x^2-91x+42x-49=013x(6x-7)+7(6x-7)=06x-7=0,x=7/6now x=1+r/100=7/6r/100=7/6-1=1/6r=(1/6)*100=50/3=16(2/3)...